OpenStudy (anonymous):

Plz Help! 1) Simplify the following algebraic expressions. -6x + 5 + 12x -6 2(x - 9) + 6(-x + 2) + 4x 3x2 + 12 + 9x - 20 + 6x2 - x (x + 2)(x + 4) + (x + 5)(-x - 1) 2) Find any real solutions for the following quadratic equations. 2 x2 - 8 = 0 x2 = -5

OpenStudy (16shuston):

For the 1st part combine all alike terms. For example 1 -6x+5+12x-6 would simplify to 6x-1

OpenStudy (anonymous):

-6x + 5 + 12x -6 add the number together and the variables together

OpenStudy (anonymous):

3x2 + 12 + 9x - 20 + 6x2 - x add the numbers together and the variables together but look to the power here the have to have the same power

OpenStudy (anonymous):

2(x - 9) + 6(-x + 2) + 4x a(c-d)=a*c-a*d use this for both then just add them

OpenStudy (anonymous):

oh ok i see

OpenStudy (anonymous):

#2 2 x2 - 8 = 0 1- add +8 for both sides 2x^2-8+8=8 2x^2=8 2- divied both sides by 2 x^2=8/2=4 3- take the sqrt of it x=sqrt(4)=+-2

OpenStudy (anonymous):

okay good luck.... Note sqrt(-#) is not a real

OpenStudy (anonymous):

thanks